Hi!
I'm not sure what exactly I did wrong in this problem. I came up with the answer 16/3 but it says that its wrong. Can you help me?
Express the statement as an equation. Use the given information to find the constant of proportionality.
M varies directly as x and inversely as y. If x = 4 and y = 3, then M = -4.
thanks!

Answer by Alan3354(69443)

Express the statement as an equation. Use the given information to find the constant of proportionality.
M varies directly as x and inversely as y. If x = 4 and y = 3, then M = -4.
----------------
M = kx/y
-4 = k*4/3
-12 = 4k
k = -3